# Interpretable and Equation-Free Response Theory for Complex Systems

### v2

@@ IV.1.3 Nonlinear Response
Using the formalism developed here, we derive easily implementable and easily interpretable formulas for the second (see Sect. II.2) and well as the arbitrary order nonlinear response (see App. A), whereby the way different intrisic time scales of the system and the corresponding modes interact with each other and with the forcing is extremely clear. Hence, as a final step proof-of-concept analysis of our system we have computed the second order Green’s functions for our system.
Indeed, we make things slightly more complicated that what has been presented in Sect. II.2. We consider the case where both forcings described above are applied. It is easy to derive that the second-order response can be written as:
@@
describes the combined effect of first applying the perturbation described by $m_{j}$ and then of the perturbation described by $m_{i}$. One should note that in general $\mathcal{G}^{(2)}_{m_{i},m_{j},\Psi}\neq\mathcal{G}^{(2)}_{m_{j},m_{i},\Psi}$ if $m_{j}\neq m_{i}$, because the time ordering matters. In our case, the second order response depends in general on all of these four terms. Given the symmetry properties of the system, if we choose $x$ as observable $\mathcal{G}^{(2)}_{m_{x},m_{x},x}=\mathcal{G}^{(2)}_{m_{y},m_{y},x}=0$, whilst $\mathcal{G}^{(2)}_{m_{x},m_{y},x}$ and $\mathcal{G}^{(2)}_{m_{y},m_{x},x}$ are in general non-vanishing. This implies that if we do not apply forcings in both direction, the second order response of $x$ vanishes. Instead, if we choose $y$ as observable, we have $\mathcal{G}^{(2)}_{m_{x},m_{y},y}=\mathcal{G}^{(2)}_{m_{y},m_{x},y}=0$, whilst $\mathcal{G}^{(2)}_{m_{x},m_{x},y}$ and $\mathcal{G}^{(2)}_{m_{y},m_{y},y}=0$ are non-vanishing. The non-vanishing second-order Green’s functions computed for RMM are reported in Fig. 4. As we see, they follow closely the functional dependence derived in Eq. 19, where in this case we obtain simple monomials because only one of the factors $\alpha_{ij}=\langle m_{p}^{T}\Pi_{j}m_{q}\Pi_{i}\Psi,\nu_{inv}\rangle$ is non-vanising, because of symmetry, for a choice $p,q=x,y$; $i,j=2,3$; and $\Psi=x,y$. It is extremely encouraging to observe that, just as for the linear case, if we repeat the same analysis using the high-resolution Ulam discretization the results are basically unchanged, compare Figs. 4-5 despite the presence of hundreds of Koopman modes, and hence of hundreds of thousands of $\alpha_{ij}$ factors, only one monomial appears to contribute to the second order Green’s function, thus supporting the efficiency of the model reduction attained with the RMM.
@@
Figure 4: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two eigenvalues of the Koopman operator. Results obtained using RMM. See Eqs. IV.1.3-30.
Figure 5: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two subdominant eigenvalues of the Koopman operator. Results obtained using the Ulam discretization. See Eqs. IV.1.3-30.
V Conclusions
@@
Constructing accurate and efficient response operators for complex systems is a problem of both theoretical and practical relevance across multiple fields in quantitative sciences [6, 2, 9, 65]. In the case of systems possessing smooth invariant measures, it is possible to resort to non-standard formulations of the FDT to recover such response operators. By combining such formalism with Koopmanism [19], one gains the important property of interpretability, as it is possible to decompose the Green’s functions of interest into a sum of terms, each associated with a specific mode of variability of the system [17, 24]. The use of Koopmanism is instrumental for establishing response formulas valid also in the case the stochastic component of a system includes jump processes [25].
@@
It is indeed possible to devise workarounds to derive Green’s functions even in the case of extremely high dimensional systems by performing suitably defined perturbation experiments, as done in the case of climate models, where the direct evaluation of response operators seems an insurmountable task [71, 72, 73]. The flip side is that one can indeed construct useful and accurate black-box-like objects that translate forcings into predicted response, but lack ability to disentangle and possibly organize hierarchically the impact of the multiple ongoing physical processes. Hence, the level of interpretability of the response operators, despite their skills, can again be disappointing.
V.1 Towards Equation-Free and Interpretable Response Formulas
Here we have developed a rather comprehensive set of response formulas - linear and nonlinear, for observables and for correlations, for static as well as for general time dependent perturbations - for Markov chains possessing a unique ergodic invariant measure, which are of primary importance for dynamical systems theory and statistical mechanics as a whole. See Appendix A and Appendix B for the general formulas. Appendix C presents a linear response formulas for the total entropy production of a Markov chain undergoing time-dependent perturbation. Hopefully these results can be useful for advancing our understanding of the sensitivity of Markov chains to perturbations and, in particular, of the response near criticality, associated with the closure of the spectral gap of the unperturbed transfer operator [41, 42]. We will focus on this specific and extremely important problem in a separate study. As for future investigations, it is also tempting to explore the case of absorbing Markov chains, which describe processes where there is a hole (or a trap) in the reference state space, so that the process is eventually killed [74]. In many cases it is possible to define and prove the existence of quasi-stationary and quasi-ergodic measures, which are constructed by adapting the usual notions to this specific case where one needs to take into account of the continuous leaks occurring in the state space [75]. Since the existence of such measures require, roughly speaking, the presence of a spectral gap of the transfer operator, it seems interesting to explore whether response formulas could be developed also for such Markov chains.
@@ V.1 Towards Equation-Free and Interpretable Response Formulas
    can be cast as simple algebraic operations performed with matrices, thus taking full advantage of the outstanding development occurred in the last decades in numeical linear algebra and the vast availability of dedicated software environments;
@@
We have thoroughly explored the efficacy of our strategy on a simple 2D gradient flow forced by additive and diagonal gaussian noise. The system we have investigated features three competing metastable states. We have been able to construct linear and nonlinear response operators that have allow us to define sensitivity and explore response to time-dependent perturbations for observables as well as for correlations of the system. We have shown that even considering a very severely reduced discrete representation of the system, we are able obtain high-quality information on its response to perturbations and to associated the response to specific modes of unperturbed variability.
@@
The framework we have in mind foresees the use of suitable methods of reduction of complexity of a system via MSM before the response theory developed here is used. MSM is very effective in creating a surrogate representation of a possibly multiscale, many particle system in a moderate number of states [39]. All one needs is a reference dataset plus few extra datasets produced with a slightly perturbed dynamics, thus allowing to obtain an estimate of the unperturbed Perron-Frobenius operator and of its perturbation in the basis defined by the MSM. Indeed, one can then study the response of the system directly at the desired coarse grained level, bypassing the need to look into all the intricacies of the system in the original resolution. The wide availability of software tools facilitating the construction of MSM [76, 77, 78], the wide range of areas of applications for MSM [39, 37], recently extended to also to climate applications [79], and the growing evidence of the efficacy of MSM in capturing the correlation properties of the full system [80] supports the strategy proposed here.
Markov State Modelling often relies on constructing the finite state space by taking advantage of optimal Voronoï tessellations of the phase space of the system. Hence, we see a clear link between the pipeline discussed above and the combination of response theory and Koopmanism - see our recent attempt in [24] - where the operation of constructing the approximate Koopman operator is performed using the Multiplicative DMD [26], which, in turn, relies on using the characteristic functions of the cells of a Voronoï tessellation as Koopman dictionary. The relationship between these two reduced order modelling approaches will be explored in the future.
V.2 A Comment of the Prony Method
@@
A final comment follows. In [55, 81] we explored the fact that response theory provides solid foundations for a key statistical analysis method used in climate science, the optimal fingerprinting method for detection and attribution of climate change [82, 83, 84, 85]. Also the results presented in this contribution seem to provide some clarification to a widely used statistical analysis method used, in this case, in many signal processing applications. In Eqs. 7-8 we have shown that the Koopman operator-based expansion of the linear Green’s functions allows to decomposite it in a sum of exponential terms with weighting factors that depend on the chosen observables and on the applied forcings, whilst the exponential decay rates depend exclusively on the properties of the unperturbed system. Indeed, this functional representation points directly to the popular Prony method, which aims at representing the multivariate response collected at discrete times of a general system to instantaneous perturbations as a weighted sum of exponentials [86, 87, 88, 89]. Usually, the number of exponentials one needs to use is a free and uncertain metaparameter of the statistical method. Our approach provides an interpretation of such a metaparameter, which corresponds to the number of discrete states (plus one) we consider in a hypothetical Markov state representation of the system. The well-known uncertainty in defining the optimal value for the metaparameter in the presence of strong noise and/or limited amount of available data can be linked to the difficulty in constructing an accurate Markov model in such conditions.
Acknowledgements
@@
This research has been partially supported by the Horizon Europe Project ClimTIP (Grant No. 100018693) and by the EPSRC project LINK (Grant No. EP/Y026675/1).
@@ Appendix B Dynamic Response of Correlations
We extend here the results presented in Sect. III to the more general case $\mathcal{M}\rightarrow\mathcal{M}_{\epsilon,n}=\mathcal{M}+\epsilon f(n)m$. Since the system has explicit time-dependent dynamics, so that the statistical properties are indexed by the observation time $n$. We have:

### v3

@@ 4.1.3 Nonlinear Response
Using the formalism developed here, we derive easily implementable and easily interpretable formulas for the second (see Subsect. 22.2) and well as the arbitrary order nonlinear response (see App. A), whereby the way different intrinsic time scales of the system and the corresponding modes interact with each other and with the forcing is extremely clear. Hence, as a final step proof-of-concept analysis of our system we have computed the second order Green’s functions for our system.
Indeed, we consider here a slightly more complicated setting than what has been earlier presented in Subsect. 22.2. We consider the case where both forcings described above are applied. It is easy to derive that the second-order response can be written as:
@@
describes the combined effect of first applying the perturbation described by $m_{j}$ and then of the perturbation described by $m_{i}$. One should note that in general $\mathcal{G}^{(2)}_{m_{i},m_{j},\Psi}\neq\mathcal{G}^{(2)}_{m_{j},m_{i},\Psi}$ if $m_{j}\neq m_{i}$, because the time ordering matters. In our case, the second order response depends in general on all of these four terms.
Given the symmetry properties of the system, if we choose $x$ as observable $\mathcal{G}^{(2)}_{m_{x},m_{x},x}=\mathcal{G}^{(2)}_{m_{y},m_{y},x}=0$, whilst $\mathcal{G}^{(2)}_{m_{x},m_{y},x}$ and $\mathcal{G}^{(2)}_{m_{y},m_{x},x}$ are in general non-vanishing. This implies that if we do not apply forcings in both direction, the second order response of $x$ vanishes. Instead, if we choose $y$ as observable, we have $\mathcal{G}^{(2)}_{m_{x},m_{y},y}=\mathcal{G}^{(2)}_{m_{y},m_{x},y}=0$, whilst $\mathcal{G}^{(2)}_{m_{x},m_{x},y}$ and $\mathcal{G}^{(2)}_{m_{y},m_{y},y}=0$ are non-vanishing.
@@
The non-vanishing second-order Green’s functions computed for RMM are reported in Fig. 4. As we see, they follow closely the functional dependence derived in Eq. 19, where in this case we obtain simple monomials because only one of the factors $\alpha_{ij}=\langle m_{p}^{T}\Pi_{j}m_{q}\Pi_{i}\Psi,\nu_{inv}\rangle$ is non-vanising, because of symmetry, for a choice $p,q=x,y$; $i,j=2,3$; and $\Psi=x,y$. It is extremely encouraging to observe that, just as for the linear case, if we repeat the same analysis using the high-resolution Ulam discretization the results are basically unchanged, compare Figs. 4-5 despite the presence of hundreds of Koopman modes, and hence of hundreds of thousands of $\alpha_{ij}$ factors, only one monomial appears to contribute to the second order Green’s function, thus supporting the efficiency of the model reduction attained with the RMM.
@@
Figure 4: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two eigenvalues of the Koopman operator. Results obtained using RMM. See Eqs. 4.1.3-30.
Figure 5: Same as in Fig. 4, but here the the results have been obtained using the Ulam discretization.
5 Conclusions
@@
Constructing accurate and efficient response operators for complex systems is a problem of both theoretical and practical relevance across multiple fields in quantitative sciences [6, 2, 9, 89]. In the case of systems possessing smooth invariant measures, it is possible to resort to non-standard formulations of the FDT to recover such response operators. By combining such a formalism with Koopmanism [20], one gains the important property of interpretability, as it is possible to decompose the Green’s functions of interest into a sum of terms, each associated with a specific mode of variability of the system [17, 25]. The use of Koopmanism is instrumental for establishing response formulas valid also in the case the stochastic component of a system includes jump processes [27].
@@
It is indeed possible to devise workarounds to derive Green’s functions even in the case of extremely high dimensional systems by performing suitably defined perturbation experiments, as done in the case of climate models, where the direct evaluation of response operators seems an insurmountable task [95, 96, 97]. The flip side is that one can indeed construct useful and accurate black-box-like objects that translate forcings into predicted response, but lack the ability to disentangle and possibly organise hierarchically the impact of the multiple ongoing physical processes. Hence, the level of interpretability of the response operators, despite their skills, can again be disappointing.
5.1 Towards Equation-Free and Interpretable Response Formulas
Here we have developed a rather comprehensive set of response formulas - linear and nonlinear, for observables and for correlations, for static as well as for general time dependent perturbations - for Markov chains possessing a unique ergodic invariant measure, which are of primary importance for dynamical systems theory and statistical mechanics as a whole. See Appendix A and Appendix B for the general formulas. Appendix C presents a linear response formulas for the total entropy production of a Markov chain undergoing time-dependent perturbation. All of the matrix equations reported in this paper are pseudo-codes that can be seamlessly translated into functioning routines using e.g. software like MATLAB [98] (used here by the author) or open source equivalents like Julia [99], Python [100], and Octave [101], among others.
Hopefully these results can be useful for advancing our understanding of the sensitivity of Markov chains to perturbations and, in particular, of the response near criticality, associated with the closure of the spectral gap of the unperturbed transfer operator [59, 102, 103, 104, 16, 17]. We will focus on this specific and extremely important problem in a separate study. As for future investigations, it is also tempting to explore the case of absorbing Markov chains, which describe processes where there is a hole (or a trap) in the reference state space, so that the process is eventually killed [105]. In many cases it is possible to define and prove the existence of quasi-stationary and quasi-ergodic measures, which are constructed by adapting the usual notions to this specific case where one needs to take into account of the continuous leak of mass occurring in the state space [106]. Since the existence of such measures requires, roughly speaking, the presence of a spectral gap of the transfer operator, it seems interesting to explore whether response formulas could be developed also for such Markov chains.
@@ 5.1 Towards Equation-Free and Interpretable Response Formulas
    can be cast as simple algebraic operations performed with matrices, thus taking full advantage of the outstanding development occurred in the last decades in numerical linear algebra and the vast availability of dedicated software environments;
@@
We have thoroughly explored the efficacy of our strategy on a simple 2D gradient flow forced by additive and diagonal gaussian noise. The system we have investigated features three competing metastable states. We have been able to construct linear and nonlinear response operators that have allow us to define sensitivity and explore response to time-dependent perturbations for observables as well as for correlations of the system. We have shown that even considering a very severely reduced discrete representation of the system, we are able to obtain high-quality information on its response to perturbations and to associated the response to specific modes of unperturbed variability.
@@
The framework above foresees the use of suitable methods of reduction of complexity of a system via MSM before the response theory developed here is used. MSM is very effective in creating a surrogate representation of a possibly multiscale, many-particle system in a moderate number of states [42]. What we propose here is to extend MSM in such a way that fairly general perturbations can be dealt with. All one needs is a reference dataset plus few extra datasets produced with a slightly perturbed dynamics, thus allowing to obtain an estimate of the unperturbed Perron-Frobenius operator and of its perturbation in the basis defined by the MSM. Indeed, one can then study the response of the system directly at the desired coarse grained level, bypassing the need to look into all the intricacies of the system in the original resolution. The wide availability of software tools facilitating the construction of MSM [107, 108, 109], the wide range of areas of applications for MSM [42, 40], recently extended to also to climate applications [110], and the growing evidence of the efficacy of MSM in capturing the correlation properties of the full system [111] support our strategy.
As mentioned above, constructing a discrete-time, finite-state Markov chain representation of a system amounts to applying the Koopman formalism where the elements of the dictionary are the characteristic functions of the Voronoï cells of the phase space. Hence, the results presented here can be seen as linking our attempts at combining of response theory and Koopmanism [17, 25, 26] with the multiplicative DMD algorithm [58]. The multiplicative DMD could be improved by using bisecting k-means to create the centroids of the Voronoï tessellation, as this methodology is better suited for detecting multiscale features and adapts well to high-dimensional datasets [112]. Increased accuracy in the estimate of response operators via Koopman methods is likely to be obtained by adopting more sophisticated DMD methods that avoid the problem of spectral pollution [113].
Concluding, it is worth noting that in many real-life applications, markovianity is a convenient, useful lie and the lack thereof is the honest, uncomfortable truth. Hence, in the future efforts should be directed at extending the results presented here to the case of Markov chains with memory [114] space and hidden Markov models [115].
5.2 A Final Comment on the Prony Method
@@
In [77, 116] we explored the fact that response theory provides solid foundations for a key statistical analysis method used in climate science, the optimal fingerprinting method for detection and attribution of climate change [117, 118, 119, 120]. Also the results presented in this contribution seem to provide some clarification for a widely used statistical analysis method used, in this case, in many signal processing applications. In Eqs. 7-8 we have shown that the Koopman operator-based expansion of the linear Green’s functions allows to decomposite it in a sum of exponential terms with weighting factors that depend on the chosen observables and on the applied forcings, whilst the exponential decay rates depend exclusively on the properties of the unperturbed system. Indeed, this functional representation points directly to the popular Prony method, which aims at representing the (in general, multivariate) response collected at discrete times of a general system to instantaneous perturbations as a weighted sum of exponentials [69, 70, 71, 72]. Usually, the number of exponentials one needs to use is a free and uncertain metaparameter of the statistical method. Our approach provides an interpretation of such a metaparameter, which corresponds to the number of discrete states (plus one) we consider in a hypothetical Markov state representation of the system. The well-known uncertainty in defining the optimal value for the metaparameter in the presence of strong noise and/or limited amount of available data can be linked to the difficulty in constructing an accurate Markov model in such conditions. Our observation is de facto dual to the findings presented in [121], where it was shown that a multivariate version of the Prony method can be used to perform DMD analysis of a nonlinear dynamical system.
[…]
VL wishes to thank M. Colbrook, I. Mezić, A. Souza, and N. Zagli and three anonymous reviewers for having provided very insightful comments on the first version of this paper.
[…]
The data needed to reproduce the figures included in the paper and the key components of the code are available here: https://figshare.com/s/60cc002c081fdcbadabc.
[…]
The author declares no competing interests.
[…]
@@
This research has been partially supported by the Horizon Europe Project ClimTIP (Grant No. 100018693), by the ARIA SCOP-PR01-P003 - Advancing Tipping Point Early Warning AdvanTip project, by the European Space Agency project PREDICT (contract 4000146344/24/I-LR), by the EPSRC project LINK (Grant No. EP/Y026675/1), and by the CROPS RETF project funded by the University of Reading.
@@ Appendix B Dynamic Response of Correlations
We extend here the results presented in Sect. 3 to the more general case $\mathcal{M}\rightarrow\mathcal{M}_{\epsilon,n}=\mathcal{M}+\epsilon f(n)m$. Since the system has explicit time-dependent dynamics, so that the statistical properties are indexed by the observation time $n$. We have: